If x - 10 = 35, which property of equality for real numbers justifies that x = 45. show the use of this property
Answer by ptaylor(2198)


x-10=35 Add 10 to both sides of the equation
10=10
x-10+10=35+10 collect like terms
x=45
"If equals are added to equals the sums are equal"

Hope this helps ----ptaylor